# Nef cones of Hilbert schemes of points on some K3 surfaces

### v1

@@ Abstract.
Let $X$ be a Mori dream K3 surface of Picard rank $2$. We compute the nef cone of the Hilbert scheme $X^{[n]}$ of $n$ points on $X$ for large $n$ using Bridgeland stability methods. We also find bounds on the nef cone for small $n$. Studying the geometry of a 22-to-11 cover $X\rightarrow Q$ of a smooth quadric $Q$ in $\mathbb{P}^{3}$, and its induced rational map $X^{[2]}\dashrightarrow Q^{[2]}$ on the Hilbert schemes, we calculate the nef cone of $X^{[2]}$ in a particular case. Finally, we also find the nef cone of the nested Hilbert scheme $X^{[n,n+1]}$ for large $n$.
@@ 1. Introduction
Given a projective variety $Y$, the ample cone in its Néron–Severi space encodes how $Y$ embeds into projective space. Understanding this cone is a key step in running the minimal model program on $Y$. The nef cone, $\operatorname{Nef}(Y)$, is the closure of the ample cone and serves as a fundamental birational invariant. However, computing $\operatorname{Nef}(Y)$ explicitly is often difficult. One has to first prove that a given divisor is nef, and then show that the nef cone does not extend further. The latter typically requires producing curves orthogonal to the candidate boundary divisors. For the former, one often pulls back known nef divisors along suitable morphisms. A fairly recent technique involving Bridgeland stability conditions developed by Bayer and Macrì (see [BM14, Lemma 3.3]) and several other authors (see [BHL+16]) has been used in [BHL+16, Kop21] to compute the nef cones of Hilbert schemes of points on hypersurfaces in $\mathbb{P}^{3}$ and on blow-ups of $\mathbb{P}^{2}$ for example.
Hilbert schemes of points form an important class of parameter spaces, and in this article we explore the question of finding the nef cones of these Hilbert schemes on K3 surfaces. As a corollary, we also obtain some information about the nested Hilbert schemes in one case.
@@ 1.1. What is known
In [BM14], Bayer and Macrì develop a powerful tool (called the Positivity Lemma) to create nef divisor classes on moduli stacks of Bridgeland semistable objects on smooth projective varieties. They also explain when a curve class is orthogonal to such a nef divisor class. They then use this technique to compute the nef cone of the Hilbert schemes $X^{[n]}$ of $n$ points on Picard rank 1 K3 surfaces $X$ for large $n$. See [BM14, Lemma 3.3, Proposition 10.3]. Here, we apply their technique to do the Picard rank 2 case.
Based on the Positivity Lemma, a general technique for constructing (extremal) nef divisor classes on $X^{[n]}$ (for large enough $n$) is discussed in [BHL+16]. Finally, using this technique, Kopper (see [Kop21]) has studied the nef cones $X^{[n]}$ (with $n\geq 3$) for rational elliptic surfaces $X$. The main idea is to consider a certain cone $\Lambda$ consisting of divisor classes that are non-negative on some particular curves on $X^{[n]}$, so that $\operatorname{Nef}(X^{[n]})\subseteq\Lambda$. Then he bounds $\Lambda$ by a larger cone, and utilizes that to show that the extremal divisors of $\Lambda$ are actually nef (using the method from [BHL+16] and the Weyl group action on $\operatorname{NE}(X)$).
For small $n$, one usually needs to utilize more classical techniques, using the geometry of the surface in consideration in much more depth. See [Rah24] for example, where the author studies the nef cones of $X^{[n]}$ for hypersurfaces $X$ in $\mathbb{P}^{3}$, with $n$ small.
@@
1.2. Our main results
Utilizing the results and ideas referred to above, we compute the nef cones of the Hilbert schemes $X^{[n]}$ (for $n\gg 0$), where $X$ is a K3 surface of Picard rank $2$ that is also a Mori dream surface. We need to briefly introduce some notation before stating our first result. Mori dream K3 surfaces of Picard rank $2$ have divisor class group $\operatorname{Cl}(X)\cong\mathbb{Z}w_{1}\oplus\mathbb{Z}w_{2}$ for some smooth curves $w_{1}$ and $w_{2}$ on $X$ that are either rational or elliptic. We use the following helpful terminology:
@@ 1.2. Our main results
    More precisely, the above assertions hold whenever $n$ satisfies the inequalitites $n>k$ and
    $32(k^{2}+2k+1)n^{4}-4k(25k^{2}+41k+16)n^{3}+4k^{2}(26k^{2}+33k+7)n^{2}-4k^{3}(9k^{2}+8k-1)n-k^{4}\geq 0.$
@@
    More precisely, the above assertions hold whenever $n$ satisfies the inequalitites $n\geq k\geq 3$ and
    $8n^{2}+(2-9k)n+8\geq 0.$
As mentioned above, the Bridgeland stability technique works for large $n$. For small $n$ however, one often needs to invoke much more classical techniques in order to find the nef cone of $X^{[n]}$. Before performing such detailed analysis, we obtain a bound for the nef cone for small $n$ using our result mentioned above.
@@
In case I when $n=k=2$, the nef cone of the Hilbert scheme $X^{[2]}$ is spanned by $w_{1}^{[2]},w_{2}^{[2]}$ and
@@
Let $X$ be in case I. Let $D_{n}$ be the extremal ray of $\mathrm{Nef}(X^{[n]})$ which does not come from the inclusion $\mathrm{Nef}(X)\hookrightarrow\mathrm{Nef}(X^{[n]})$. Then the nef cone of the nested Hilbert scheme $X^{[n,n+1]}$ for $n\gg 0$, is spanned by
$\{\mathrm{pr}_{a}^{*}(w_{1}^{[n]}),\mathrm{pr}_{a}^{*}(w_{2}^{[n]}),w_{1}^{\mathrm{diff}},w_{2}^{\mathrm{diff}},\mathrm{pr}_{a}^{*}(D_{n}),\mathrm{pr}_{b}^{*}{D_{n+1}}\}.$
@@
In §2, we review the necessary background on K3 surfaces, Hilbert schemes, and Bridgeland stability conditions, and outline our overall strategy (see §2.5). In §3, we present the main results. §3.1 focuses on the large $n$ case, where we use Bridgeland stability to determine nef cones. Finally in §3.2 we address the small $n$ case.
@@ Acknowledgments
We would like to thank Izzet Coskun and Jack Huizenga for several helpful discussions.
@@ 2.5. Strategy:
We then show that $\Lambda$ is the cone in $N^{1}(X^{[n]})$ spanned by $\operatorname{Nef}(X^{[n]})$ and another divisor $D.$ The final step is to show that $D$ is a nef divisor using the following result,
@@ 3.1. The nef cone for large $n$
by inequality (3.1.1) above. Similary,
@@
By [BHL+16, Theorem 3.6], $W(\mathcal{O}_{X}(-w_{1}),I_{Z})$ is the Gieseker wall for $\mathbf{v}(I_{Z})$ in the $(A,P)$-slice.
@@
    More precisely, the above assertions hold whenever $n$ satisfies the inequalitites $n>k$ and
    $32(k^{2}+2k+1)n^{4}-4k(25k^{2}+41k+16)n^{3}+4k^{2}(26k^{2}+33k+7)n^{2}-4k^{3}(9k^{2}+8k-1)n-k^{4}\geq 0.$
@@
    More precisely, the above assertions hold whenever $n$ satisfies the inequalitites $n\geq k\geq 3$ and
    $8n^{2}+(2-9k)n+8\geq 0.$
@@ 3.2. Bounds for small $n$
If $a_{1}w_{1}^{[n+1]}+a_{2}w_{2}^{[n+1]}-B^{[n+1]}/2$ is nef on $X^{[n+1]}$, then $a_{1}w_{1}^{[n]}+a_{2}w_{2}^{[n]}-B^{[n]}/2$ is nef on $X^{[n]}$. This is because of the following reason: suppose $a_{1}w_{1}^{[n]}+a_{2}w_{2}^{[n]}-B^{[n]}/2$ is negative on a curve $C$ on $X^{[n]}$. Then the same curve $C$ can be viewed as a curve on $X^{[n+1]}$ by “adding” a fixed general point of $X$ to each scheme on $C$ and this does not change intersection numbers. So, $\operatorname{Nef}(X^{[n]})\supseteq\operatorname{Nef}(X^{[n+1]})$, where we identify $w_{i}^{[n]}$ with $w_{i}^{[n+1]}$ for $i=1,2$, and identify $B^{[n]}$ with $B^{[n+1]}$. We are going to use this fact in the proof below.
@@
All that remains to show is that $C$ indeed carries a $g^{1}_{k}$. Suppose $n=k\geq 3$. We recall that the Brill-Noether Theorem tells us, in particular, that if the Brill-Noether number
@@
Let $X$ be as in 3.4. For the case $n=k$, there exists an effective class $C]\in N_{1}(X^{[n]})$ dual to the class $\Gamma\coloneq w_{1}^{[n]}+w_{2}^{[n]}-\frac{1}{2}B^{[n]}$.
Finally, we want to show that when $n=k=2$, the divisor for which we produced a dual curve above is actually nef. Then the existence of the dual curve guarantees that it is extremal. Thus we get the following theorem.
@@
In case I when $n=k=2$, the nef cone of the Hilbert scheme $X^{[2]}$ is spanned by $w_{1}^{[2]},w_{2}^{[2]}$ and
@@
The remainder of this article proves this assertion.
@@
There are two classes of subschemes $\xi\subset X$ that arise as points of $X^{[2]}$: Pairs of distinct points $\{x_{1},x_{2}\}$ and a length 2 subscheme supported at a singe point $p$, which we think of as the pair $\{p,v\}$ where $v\in T_{X,p}$. Let $R$ be the ramification locus of $\varphi$. Define $\phi$ as follows:
@@
As defined, $\phi$ is a morphism when restricted to the open locus $U$ of subschemes which are not fibers of $\varphi$ and which are not supported entirely along the ramification curve $R$. Furthermore, since $\mathrm{codim}_{X}(X\setminus U)=2$, we can extend the pullback of any line bundle $L\in\mathrm{Pic}(Q^{[2]})$ to $X^{[2]}$. Applying this construction to $\phi$, we get line bundles $\phi^{*}L_{1}^{[2]},\phi^{*}L_{2}^{[2]},\phi^{*}\frac{B_{Q}}{2}\in\mathrm{Pic}(X^{[2]})$. Since both Hilbert schemes are smooth, by the construction above $\phi^{*}L_{i}^{[2]}\cong w_{i}^{[2]}$ and
@@
In order to show that $\phi^{*}(L_{1}^{[2]}+L_{2}^{[2]}-\frac{B_{Q}}{2})$ is nef, we need to examine the base locus of $\phi$. Since $\varphi^{-1}(p)=\{q_{1},q_{2}\}\subset X$ (not necessarily distinct), we have a flat family $\mathcal{X}\to Q$. This induces a rational map $\varphi^{\prime}:Q\dashrightarrow X^{[2]}$. By the universal property of the Hilbert scheme, $\varphi^{\prime}$ is a morphism, and $Q^{\prime}\coloneq\varphi^{\prime}(Q)\cong Q$. Similarly, let $R\subset X$ be the locus where $d\varphi:T_{X}\to\varphi^{*}T_{Q}$ drops rank. Note that $d\varphi$ drops rank by exactly 1 and never 2, since the condition of dropping rank is a condition on the rank of $d\varphi$, and the expected dimension of the corank 2 locus is at most 0. By Riemann-Hurwitz, $R\in|2w_{1}+2w_{2}|$ Note that points in the total space of $T_{X}|_{R}\to R$ are the length 2 subschemes of $X$ whose supports are a single point of $R$. Set $S\coloneq\mathbb{P}T_{X}|_{R}$ By the same argument as above, we get a map $(d\varphi)^{\prime}:S\to X^{[2]}$. Thus,
@@
We analyze the components independently, beginning with $S\coloneq\mathbb{P}T_{X}|_{R}$. This locus is a ruled surface over the curve $R$, so $\mathrm{Pic}(S)$ is generated by two elements $E$ and $F$, where $F$ is the class of a fiber of the ruled surface over a point of $R$ and $E$ is a section. First, observe that for $r\in R$ a closed point, the fiber $S_{r}$ corresponds to the distinct tangent directions of $X$ at $r$. It follows that $[F]$ is the class of a curve in $X^{[2]}$ which is contained entirely in $B_{X}$, the exceptional locus of the Hilbert-Chow morphism. In conclusion,
@@
Recall the theorem of Fujita (see [Fuj83, Theorem 2.8]), which says that if a line bundle $\mathcal{O}(D)$ on a projective variety restricts to an ample line bundle on the base locus $\mathrm{Bs}(|D|)$, then $\mathcal{O}(D)$ is semiample, i.e. $\mathcal{O}(mD)$ is basepoint free for $m\gg 0$. Since a semiample line bundle is nef, it is sufficient to know that the candidate nef class
@@
Future directions
There are several ways in which the preceding may be generalized. Most directly, the number of points in the Hilbert scheme in the preceding proof can be increased. Carrying out similar, albeit harder analysis for the base locus of the rational map to $Q^{[n]}$ would then yield an analogous result for those larger numbers of points. Another direction is to use this technique to compute for K3s which are finite covers of other surfaces for which nef cone information is already known. The most difficult direction seems to be the nested Hilbert schemes, as those will in general fail to be smooth.

### v2

@@ Abstract.
We illustrate the typical usage of Bayer and Macrì’s Positivity Lemma to compute the nef cones of the Hilbert schemes $X^{[n]}$ by combining the Bridgeland stability methods (for large $n$) and classical methods (for small $n$). We use Mori dream K3 surfaces $X$ of Picard rank $2$ as our working example. We also compute the nef cones of the nested Hilbert schemes $X^{[n,n+1]}$ for such $X$, for large $n$.
@@ 1. Introduction
Given a projective variety $Y$, the ample cone in its Néron–Severi space encodes how $Y$ embeds into projective space. Understanding this cone is a key step in running the minimal model program on $Y$. The nef cone, $\operatorname{Nef}(Y)$, is the closure of the ample cone and serves as a fundamental birational invariant. However, computing $\operatorname{Nef}(Y)$ explicitly is often difficult. One has to first prove that a given divisor is nef, and then show that the nef cone does not extend further. The latter typically requires producing curves orthogonal to the candidate boundary divisors. For the former, one often pulls back known nef divisors along suitable morphisms. A fairly recent technique involving Bridgeland stability conditions developed by Bayer and Macrì (see [BM14b, Lemma 3.3]) and several other authors (see [BHL+16]) has been used in [BHL+16, Kop21] to compute the nef cones of Hilbert schemes of points on hypersurfaces in $\mathbb{P}^{3}$ and on blow-ups of $\mathbb{P}^{2}$ for example.
Hilbert schemes of points form an important class of parameter spaces, and in this article we explore the question of finding the nef cones of these Hilbert schemes on K3 surfaces. As a corollary, we also obtain information about the nested Hilbert schemes.
@@ 1.1. What is known
In [BM14b], Bayer and Macrì develop a powerful tool (called the Positivity Lemma) to create nef divisor classes on moduli stacks of Bridgeland semistable objects on smooth projective varieties. They also explain when a curve class is orthogonal to such a nef divisor class. They then use this technique to compute the nef cone of the Hilbert schemes $X^{[n]}$ of $n$ points on Picard rank 1 K3 surfaces $X$ for large $n$. See [BM14b, Lemma 3.3, Proposition 10.3]. In fact, their methods (see [BM14a, Theorem 12.1]) compute the nef cone of the Hilbert scheme $X^{[n]}$ for any $n$ for any K3 surface $X$. Bridgeland stability techniques are also utilized by Li and Zhao ([LZ19]) on $X=\mathbb{P}^{2}$, by Yanagida and Yoshioka ([YY14, Yos16]) on abelian surfaces $X$, and by Nuer and Yoshioka [Nue16],[NY20] on Enriques surfaces $X$ to answer similar questions. See also [ABCH13].
Based on the Positivity Lemma, a general technique for constructing (extremal) nef divisor classes on $X^{[n]}$ (for large enough $n$) is discussed in [BHL+16]. Using this technique, Kopper (see [Kop21]) has studied the nef cones $X^{[n]}$ (with $n\geq 3$) for rational elliptic surfaces $X$. The main idea is to consider a certain cone $\Lambda$ consisting of divisor classes that are non-negative on some particular curves on $X^{[n]}$, so that $\operatorname{Nef}(X^{[n]})\subseteq\Lambda$. Then he bounds $\Lambda$ by a larger cone, and utilizes that to show that the extremal divisors of $\Lambda$ are actually nef (using the method from [BHL+16] and the Weyl group action on $\operatorname{NE}(X)$). For small $n$, one usually needs to utilize more classical techniques, using the geometry of the surface in consideration in much more depth. See [Rah24] for example, where the author studies the nef cones of $X^{[n]}$ for hypersurfaces $X$ in $\mathbb{P}^{3}$, with $n$ small.
For K3 surfaces, however, several favorable properties hold, allowing one to eliminate the “large enough nn” requirement. In particular, [BM14a, Theorem 5.7] provides a detailed description of the behavior of walls and divisorial contractions, many of which are governed by solutions to certain Pell’s equations. Together with [BM14a, Theorem 1.2], this makes it possible to compute the movable cone of $X^{[n]}$ for a K3 surface $X$. Moreover, [BM14a, Theorems 12.1 and 12.3] allow one to compare the nef and the movable cones of $X^{[n]}$.
@@
1.2. What we do in this paper
Our primary goal in this primarily expository article is to illustrate the typical usage of Bayer and Macrì’s Positivity Lemma and the methods developed in [BHL+16] to compute the nef cones of the Hilbert schemes $X^{[n]}$ by combining the Bridgeland stability methods (for large $n$) and classical methods (for small $n$). We use K3 surfaces $X$ of Picard rank $2$ that are also Mori dream surfaces as our working example. However, we would like to remind the reader that for K3 surfaces, one can simply utilize the strategy mentioned earlier (more specifically, [BM14a, Theorem 12.1]) to get the desired result for any $n$.
We need to briefly introduce some notation before stating the first result. Mori dream K3 surfaces of Picard rank $2$ have divisor class group $\operatorname{Cl}(X)\cong\mathbb{Z}w_{1}\oplus\mathbb{Z}w_{2}$ for some smooth curves $w_{1}$ and $w_{2}$ on $X$ that are either rational or elliptic. We use the following helpful terminology:
@@ 1.2. What we do in this paper
    More precisely, the above assertions hold whenever $n$ satisfies the inequalities $n>k$ and 32(k^2+2k+1)n^4-4k(25k^2+41k+16)n^3+4k^2(26k^2+33k+7)n^2-4k^3(9k^2+8k-1)n-k^4≥0.
@@
    More precisely, the above assertions hold whenever $n$ satisfies the inequalities $n\geq k\geq 3$ and 8n^2+(2-9k)n+8≥0.
The above calculation can now be used to obtain a bound for the nef cone of $X^{[n]}$ for small $n$.
@@
In case I when $k=2$, the nef cone of the Hilbert scheme $X^{[2]}$ is spanned by $w_{1}^{[2]},w_{2}^{[2]}$ and
@@
In subsection 3.3, we once again calculate the nef cone of the Hilbert scheme $X^{[2]}$ in case I when $k=2$, this time directly using [BM14a, Theorem 12.1]. In other words, we bypass [BHL+16] and illustrate a direct computation of the nef cone of $X^{[n]}$ in the case of K3 surfaces $X$.
@@
Let $X$ be in case I. Let $D_{n}$ be the extremal ray of $\mathrm{Nef}(X^{[n]})$ which does not come from the inclusion $\mathrm{Nef}(X)\hookrightarrow\mathrm{Nef}(X^{[n]})$. Then the nef cone of the nested Hilbert scheme $X^{[n,n+1]}$ for $n\geq 9k/8$ is spanned by
$\{\mathrm{pr}_{a}^{*}(w_{1}^{[n]}),\mathrm{pr}_{a}^{*}(w_{2}^{[n]}),w_{1}^{\mathrm{diff}},w_{2}^{\mathrm{diff}},\mathrm{pr}_{a}^{*}(D_{n}),\mathrm{pr}_{b}^{*}(D_{n+1})\}.$
In fact, using the complete description (see [BM14a, Theorem 12.1]) of the nef cone of $X^{[n]}$ for any $n$, where $X$ is a K3 surface, one can similarly compute the nef cone of $X^{[n,n+1]}$ for any $n$.
@@
In §2, we review the necessary background on K3 surfaces, Hilbert schemes, and Bridgeland stability conditions, and outline our overall strategy (see §2.5). In §3, we present the main computations. §3.1 focuses on the large $n$ case, where we use techniques from [BHL+16] to determine nef cones. Finally in §3.2 and §3.3 we address the small $n$ case and directly apply [BM14a].
@@ Acknowledgments
We would like to thank Izzet Coskun, Jack Huizenga, and Howard Nuer for several helpful discussions. We also express our gratitude toward Arend Bayer and Emanuele Macrì for their comments on an earlier draft of this article that drew our attention to the results in [BM14a], and made the exposition clearer and more historically accurate.
@@ 2.5. Strategy:
We then show that $\Lambda$ is the cone in $N^{1}(X^{[n]})$ spanned by $\operatorname{Nef}(X^{[n]})$ and another divisor $D.$ The final step is to show that $D$ is a nef divisor using the following result.
@@ 3.1. The nef cone for large $n$
by inequality (3.1.1) above. Similarly,
@@
By [BHL+16, Theorem 3.6], $W(\mathcal{O}_{X}(-w_{1}),I_{Z})$ is the Gieseker wall for $\mathbf{v}(I_{Z})$ in the $(H,D)$-slice.
@@
    More precisely, the above assertions hold whenever $n$ satisfies the inequalities $n>k$ and 32(k^2+2k+1)n^4-4k(25k^2+41k+16)n^3+4k^2(26k^2+33k+7)n^2-4k^3(9k^2+8k-1)n-k^4≥0.
@@
    More precisely, the above assertions hold whenever $n$ satisfies the inequalities $n\geq k\geq 3$ and 8n^2+(2-9k)n+8≥0.
@@ 3.2. Bounds for small $n$
If $a_{1}w_{1}^{[n+1]}+a_{2}w_{2}^{[n+1]}-\frac{B^{[n+1]}}{2}$ is nef on $X^{[n+1]}$, then $a_{1}w_{1}^{[n]}+a_{2}w_{2}^{[n]}-\frac{B^{[n]}}{2}$ is nef on $X^{[n]}$. This is because of the following reason: suppose $a_{1}w_{1}^{[n]}+a_{2}w_{2}^{[n]}-B^{[n]}/2$ is negative on a curve $C$ on $X^{[n]}$. Then the same curve $C$ can be viewed as a curve on $X^{[n+1]}$ by “adding” a fixed general point of $X$ to each scheme on $C$ and this does not change intersection numbers. So, $\operatorname{Nef}(X^{[n]})\supseteq\operatorname{Nef}(X^{[n+1]})$, where we identify $w_{i}^{[n]}$ with $w_{i}^{[n+1]}$ for $i=1,2$, and identify $B^{[n]}$ with $B^{[n+1]}$. We are going to use this fact in the proof below.
@@
Remark 3.8.
The following shows that even if we did not bound $\operatorname{Nef}(X^{[n]})$ by two closed cones (and later show that it coincides with one of them), we can directly (in this specific case) explicitly show that the desired ray is extremal.
It is enough to show that $C$ carries a $g^{1}_{k}$. Suppose $n=k\geq 3$. We recall that the Brill-Noether Theorem tells us, in particular, that if the Brill-Noether number
@@
Let $X$ be as in 3.4. For the case $n=k$, there exists an effective class $[C]$ in $N_{1}(X^{[n]})$ dual to the class $\Gamma\coloneq w_{1}^{[n]}+w_{2}^{[n]}-\frac{1}{2}B^{[n]}$.
Finally, we want to show that when $n=k=2$, the divisor for which we produced a dual curve above is actually nef. At that point, $\operatorname{Nef}(X^{[2]})$ must be exactly the cone $\Lambda$ in 3.7, so since $\Lambda$ was a bound, the ray generated by this divisor must be extremal. The existence of the dual curve also guarantees that it is extremal. Thus we get the following theorem:
@@
In case I when $k=2$, the nef cone of the Hilbert scheme $X^{[2]}$ is spanned by $w_{1}^{[2]},w_{2}^{[2]}$ and
@@
The remainder of this subsection proves this assertion.
@@
There are two classes of subschemes $\xi\subset X$ that arise as points of $X^{[2]}$: Pairs of distinct points $\{x_{1},x_{2}\}$ and a length 2 subscheme supported at a single point $p$, which we think of as the pair $\{p,v\}$ where $v\in T_{X,p}$. Let $R$ be the ramification locus of $\varphi$. Define $\phi$ as follows:
@@
As defined, $\phi$ is a morphism when restricted to the open locus $U$ of subschemes which are not fibers of $\varphi$ and which are not supported entirely along the ramification curve $R$. Furthermore, since everything in sight is normal and $\mathrm{codim}_{X}(X\setminus U)=2$, we can extend the pullback of any line bundle $L\in\mathrm{Pic}(Q^{[2]})$ from $U$ to all of $X^{[2]}$. Applying this construction to $\phi$, we get line bundles $\phi^{*}L_{1}^{[2]},\phi^{*}L_{2}^{[2]},\phi^{*}\frac{B_{Q}}{2}\in\mathrm{Pic}(X^{[2]})$. Since both Hilbert schemes are smooth, by the construction above $\phi^{*}L_{i}^{[2]}\cong w_{i}^{[2]}$ and
@@
In order to show that $\phi^{*}(L_{1}^{[2]}+L_{2}^{[2]}-\frac{B_{Q}}{2})$ is nef, we need to examine the base locus of $\phi$. Since $\varphi^{-1}(p)=\{q_{1},q_{2}\}\subset X$ (not necessarily distinct), we have a flat family $\mathcal{X}\to Q$. This induces a rational map $\varphi^{\prime}:Q\dashrightarrow X^{[2]}$. By the universal property of the Hilbert scheme, $\varphi^{\prime}$ is a morphism, and $Q^{\prime}\coloneq\varphi^{\prime}(Q)\cong Q$. Similarly, let $R\subset X$ be the locus where $d\varphi:T_{X}\to\varphi^{*}T_{Q}$ drops rank. Note that $d\varphi$ drops rank by exactly 1 and never 2, since the condition of dropping rank is a condition on the rank of $d\varphi$, and the expected dimension of the corank 2 locus is at most 0. By Riemann-Hurwitz, $R\in|2w_{1}+2w_{2}|$ Note that points in the total space of $T_{X}|_{R}\to R$ are identified with the length 2 subschemes of $X$ whose supports are a single point of $R$. Set $S\coloneq\mathbb{P}T_{X}|_{R}$. By the same argument as above, we get a map $(d\varphi)^{\prime}:S\to X^{[2]}$. Thus,
@@
We analyze the components independently, beginning with $S$. This locus is a ruled surface over the curve $R$, so $\mathrm{Pic}(S)$ is generated by two elements $E$ and $F$, where $F$ is the class of a fiber of the ruled surface over a point of $R$ and $E$ is a section. First, observe that for $r\in R$ a closed point, the fiber $S_{r}$ corresponds to the distinct tangent directions of $X$ at $r$. It follows that $[F]$ is the class of a curve in $X^{[2]}$ which is contained entirely in $B_{X}$, the exceptional locus of the Hilbert-Chow morphism. In conclusion,
@@
Recall the theorem of Fujita (see [Fuj83, Theorem 2.8]), which says that if a line bundle $\mathcal{O}(D)$ on a projective variety restricts to an ample line bundle on the base locus $\mathrm{Bs}(|D|)$, then $\mathcal{O}(D)$ is semiample, i.e. $\mathcal{O}(mD)$ is basepoint free for $m\gg 0$. Since a semiample line bundle generates a ray of the nef cone by definition, it is sufficient to know that the candidate nef class
@@
Remark 3.17.
It is interesting to note that the extremality of the final ray shown above is already suggested by the above proof that the class is nef. We showed that the class was nef by exhibiting it as a limit. Indeed, it is exactly the border case in the Fujita theorem: its restriction to $\mathrm{Bs}(\phi)$ actually fails to be ample (via the intersection numbers computed above), but is the limit of a sequence whose elements are. It would be interesting to examine the relationship between the boundary cases in this application of Fujita’s theorem and extremality of the associated classes on the Hilbert scheme.
3.3. A direct computation of the nef cone of $X^{[2]}$
In this subsection, we use [BM14a, Theorem 12.1] to illustrate an alternate way of calculating the nef cone of $X^{[2]}$ in case I, assuming $k=2$.
Let us first recall some notation. Let $H^{*}_{\operatorname{alg}}(X,\mathbb{Z})$ be the algebraic Mukai lattice of $X$ with Mukai pairing denoted by $\langle-,$-$\rangle$. The Mukai vector $\mathbf{v}:K(X)\to H^{*}_{\operatorname{alg}}(X,\mathbb{Z})$ is defined by
$\mathbf{v}(E)=\operatorname{ch}(E)\sqrt{\operatorname{td}(X)}=(\operatorname{ch}_{0},\operatorname{ch}_{1},\operatorname{ch}_{0}+\operatorname{ch}_{2}),$
and the Mukai pairing takes the form $\langle(r,c,s),(r^{\prime},c^{\prime},s^{\prime})\rangle=(c,c^{\prime})-rs^{\prime}-r^{\prime}s$, where (,)(-,-) is the intersection pairing on $H^{2}(X,\mathbb{Z})$. Let ${\bf v}:=(1,0,1-2)=(1,0,-1)$ be the Mukai vector corresponding to the moduli space $X^{[2]}$. In other words, $X^{[2]}$ is the moduli space $M_{\sigma}({\bf v})$ of Bridgeland semistable objects with Mukai vector $\bf v$}, where $\sigma$ is a Bridgeland stability condition lying above the Gieseker wall. Given a Mukai vector $\bf w$}, we denote $\langle{\bf w},{\bf w}\rangle$ by ${\bf w}^{2}$, and its orthogonal complement by
${\bf w}^{\perp}:=\left\{{\bf w}^{\prime}\in H^{*}_{\operatorname{alg}}(X,\mathbb{Z})\,|\,\langle{\bf w},{\bf w}^{\prime}\rangle=0\right\}.$
The closed cone of positive classes defined by the Beauville–Bogomolov quadratic form on the Néron-Severi group of the hyperkähler manifold $X^{[2]}$ is denoted by $\overline{\operatorname{Pos}}(X^{[2]})$. Finally, let $\theta:{\bf v}^{\perp}\to NS(X^{[2]})$ be the isometry of lattices obtained from the Hodge-structure isomorphism between the orthogonal complement of $\bf v$} in $H^{*}(X,\mathbb{Z})$ and $H^{2}(X^{[2]},\mathbb{Z})$ (see [Yos01]). We refer the reader to [BM14a, NY20] for more details.
By [BM14a, Theorem 12.1], the nef cone of $X^{[2]}$ is cut out in $\overline{\operatorname{Pos}}(X^{[2]})$ by all linear subspaces of the form $\theta({\bf v}^{\perp}\cap{\bf a}^{\perp})$, for all classes $\bf a$$^{\perp}$ in $H^{*}_{\operatorname{alg}}(X,\mathbb{Z})$ satisfying the inequalities ${\bf v}^{2}\geq-2$ and $0\leq\langle{\bf v},{\bf a}\rangle\leq{\bf v}^{2}/2$.
By the proof of Theorem 3.10 above, we know that all the three divisor classes $w_{1}^{[2]},w_{2}^{[2]}$ and
$w_{1}^{[2]}+w_{2}^{[2]}-\frac{1}{2}B^{[2]}$
are nef on $X^{[2]}$. Thus, to prove that the nef cone of $X^{[2]}$ is spanned by $w_{1}^{[2]},w_{2}^{[2]}$ and
$w_{1}^{[2]}+w_{2}^{[2]}-\frac{1}{2}B^{[2]}$
(i.e., that it is not larger than the cone spanned by these vectors), it suffices to show that each of the three faces of the cone spanned by these three vectors is among the linear spaces $\theta({\bf v}^{\perp}\cap{\bf a}^{\perp})$ mentioned above. Since $\theta((0,-w_{1},0))=w_{1}^{[2]}$, $\theta((0,-w_{2},0))=w_{2}^{[2]}$, and $\theta((1,0,n-1))=-\tfrac{B^{[2]}}{2}$ (see [BM14b, Remark 10.4]), it follows from some straightforward computations that the Mukai vectors ${\bf a}=(1,-w_{1},1)$, $(1,-w_{2},1)$ and $(1,0,1)$ give us the desired hyperplanes.
3.4. Some closing remarks and possible future directions
It is interesting to note that the classical flavored proof for two points on one class of K3 surfaces given above only used one K3-specific ingredient. It is obviously necessary to understand the intersection theory of the surface, but that is understood in many cases. The K3-specific ingredient was stability of the restricted tangent bundle. This indicates that a similar technique might be attempted for surfaces that are not K3s. Furthermore, even the stability of the restricted tangent bundle on the nose was not totally necessary. The full version of the theorem cited from [Ful11] uses only the slope of the first graded piece of the Harder-Narasimhan filtration. As such, it should be possible to achieve similar results for other surfaces at the expense of a possibly more difficult computation. In the case that this computation is unmanageable at present, it would still be interesting to compare $T_{X}$ and $f^{*}T_{Y}$ for any finite map $X\to Y$ of surfaces to obtain a bound on positivity. Finally, working with a larger number $n$ of points should be possible by examining different vector bundles that still have some relation to ramification. Each separate $n$ requires a different computation using this method. On the other hand, the brilliance of the methods pioneered in [BM14a], [BM14b] shows in how efficiently every value of $n$ is dealt with.